8 cos x + 5 sin x

  1. 8 cos x + 5 sin x

Grade:11

1 Answers

Vijay Mukati
8 years ago
HIint.: a cosx + b sin x = sqrt(a2+b2)cos(x-alpha), where alpha = tan -1(b/a). Thanks,